OpenStudy (anonymous):

Solve by elimination 3x + 6y = 6 2x - 3y = 4 i got (2,-2)

OpenStudy (anonymous):

u here already

OpenStudy (bahrom7893):

3*2+6*(-2) = 6 - 12 =-6, not 6 the answer's wrong, try again.

OpenStudy (anonymous):

kk

OpenStudy (bahrom7893):

ori please dont give away the answer

OpenStudy (anonymous):

no im doing plz dont

OpenStudy (anonymous):

wait im stuck know

OpenStudy (bahrom7893):

post your work so far.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):



OpenStudy (bahrom7893):

x+2y=2 x = 2-2y 2(2-2y)-3y=4 4-4y-3y=4 -7y=0 y = 0

OpenStudy (anonymous):

o lol

OpenStudy (bahrom7893):

what's x?

OpenStudy (anonymous):

let me plug in y right

OpenStudy (anonymous):

with 0

OpenStudy (bahrom7893):

yup

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

x=2 that what i had b4 am i wrong again

OpenStudy (bahrom7893):

Let's check: 3x + 6y = 6 2x - 3y = 4 You have: (2, 0) 3*2+6*0 = 6 + 0 = 6 (correct) 2*2 - 3*0 = 4 - 0 = 4 (correct)

OpenStudy (anonymous):

thank you ready for another

OpenStudy (bahrom7893):

go for it

OpenStudy (anonymous):

thx